Sales total $300,000 when variable costs total $180,000 and fixed costs are $60,000. Breakeven sales total:

            A.        $300,000

            B.        $150,000

            C.        $ 60,000

            D.        $ 90,000

            E.         Cannot be determined with information provided.
Answer #1

Contribution margin=Sales-Variable cost

=(300,000-180,000)=$120,000

Contribution margin ratio=Contribution margin/Sale

=120,000/300,000=0.4

Hence breakeven sales=Fixed cost/Contribution margin ratio

=60,000/0.4

=$150,000